[figure: 9.7 cm]
Magnetism of T h(F eAl)_{12} alloys

K. Rećko^{a}, L. Dobrzyski^{a}^{b}, K. Szymaski^{a}, D. Satua^{a}, K. Perzyska^{a}, M. Biernacka^{a}, P. Zaleski^{a}, J. Waliszewski^{a}, W. Suski^{c}, K. Wochowski^{c}, M.

Hofmann^{d}, D. Hohlwein^{d}

aInstitute of Experimental Physics University of Biaystok, Lipowa 41, 15-424 Biaystok,

Poland bSotan Institute for Nuclear Studies, 05-400 Otwock-wierk, Poland

cInstitute of Low Temperatures and Structure Research PAN, PO Box 1410, 50-950

Wrocaw 2, Poland

dHahn-Meitner-Institut, BENSC, Glienicker Strae 100, D-14109 Berlin, Germany

We have studied the magnetic ordering of T hF e_{x}Al_{12−x} powder samples with iron concentration x = 4, 4.5 and 5 by means of x-ray and neutron diffraction techniques as well as spectroscopy with the monochromatic circularly polarized Mssbauer source (MCPMS) and magnetization measurements. Aforementioned thorium compounds crystallize in bct (I4/mmm) type structure. The magnetization measurements of T hF e_{x}Al_{12−x} family disclose the methamagnetic states. The appropriate magnetic phase transitions against increasing iron content are close to: 160, 200 and 220 K, properly. The Mssbauer investigations reveal no ferromagnetism in T hF e_{4}Al_{8} sample and distinct ferromagnetic character of T hF e_{4.5}Al_{7.5} composition. Neutron diffraction shows the long - range magnetic order in the case of T hF e_{4}Al_{8} only. The consistent description of the low temperature magnetic scattering requires the propagation wave vector in form: \overrightarrow{q} = (0.267(5), 0.267(5), 0). The magnetic moment per iron atom is 1.5(3)μ_{B} at 1.5K.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : K. Rećko

Address for correspondence : Institute of Experimental Physics, University of Bialystok, Lipowa 41, 15-424 Bialystok, Poland

Email address : karo@alpha.uwb.edu.pl